# SASSL: Enhancing Self-Supervised Learning via Neural Style Transfer

### v2

@@ 4 Proposed Method: SASSL
Recent work in self-supervised learning has shown the importance of diverse data augmentation to achieve state-of-the-art performance across downstream tasks. However, few approaches focus on transformations that preserve the semantic information of pre-training data, though this is essential for strong self-supervised representations. To tackle this limitation, we propose SASSL, a novel data-augmentation strategy based on neural style transfer that decouples semantic from appearance attributes, enforcing transformations that operate strictly on the input color and texture.
@@
Style Transfer as data preprocessing. We incorporate style transfer to the default preprocessing pipeline of SSL methods. We note that our method is not specific to a particular SSL method, and can be readily applied with different methods. Figure 1 describes an example of our SASSL data augmentation pipeline, where Style Transfer is applied after random cropping. A raw input image ${\bm{i}}_{0}$ is randomly cropped, producing a view that is taken as the content image ${\bm{i}}_{c}$. Given an arbitrary style image ${\bm{i}}_{s}$ (we discuss the choice of ${\bm{i}}_{s}$ below), the style transfer block generates a stylized image ${\bm{i}}_{\text{cs}}$ by imposing the texture attributes of ${\bm{i}}_{s}$ over ${\bm{i}}_{c}$. Finally, the stylized image ${\bm{i}}_{\text{cs}}$ is passed to the remaining data augmentation blocks to produce an augmented sample ${\bm{i}}_{\text{aug}}$.
@@
As discussed in recent work on self-supervised augmentation (Han et al., 2022; Chen et al., 2020a), adding a strong transformation to a self-supervised method tends to degrade performance. For this reason, it is crucial to control the amount of stylization imposed in the augmentation stage. We do so by introducing three hyperparameters: a stylization probability $p\in[0,1]$, which dictates whether an image is stylized or not, a blending factor $\alpha\in[0,1]$ to combine content ${\bm{z}}_{c}$ and style ${\bm{z}}_{s}$ representations, and an interpolation factor $\beta\in[0,1]$ to combine content ${\bm{i}}_{c}$ and stylized ${\bm{i}}_{\text{cs}}$ images.
@@
Given style representations extracted from content and style reference images ${\bm{z}}_{c}=\mathcal{F}({\bm{i}}_{c})$ and ${\bm{z}}_{s}=\mathcal{F}({\bm{i}}_{s})$, respectively, we obtain an intermediate stylized image $\hat{{\bm{i}}}_{\text{cs}}$ by applying a convex combination on them based on the blending factor $\alpha$.
$\hat{{\bm{i}}}_{\text{cs}}=$ $\ \mathcal{T}({\bm{i}}_{c},\hat{{\bm{z}}})$ (11)
$\hat{{\bm{z}}}=$ $\ (1-\alpha){\bm{z}}_{c}+\alpha{\bm{z}}_{s}$ (12)
@@
Then, the final stylization output is obtained via a convex combination between the intermediate stylized image $\hat{{\bm{i}}}_{cs}$ and the content image ${\bm{i}}_{c}$ based on the interpolation factor $\beta$.
@@
Algorithm 1 describes our proposed style transfer data augmentation block. Figure 3 illustrates the effect of the feature blending and image interpolation operations, showcasing their importance to control the stylization effect without degrading the semantic attributes.
Output: ${\bm{i}}_{\text{cs}}$
${\bm{z}}_{c}\leftarrow\mathcal{F}({\bm{i}}_{c})$ ;
  # Style representation of content image
${\bm{z}}_{s}\leftarrow\mathcal{F}({\bm{i}}_{s})$ ;
  # Style representation of style image
$\alpha\sim\mathcal{U}(\alpha_{\text{min}},\alpha_{\text{max}})$;
  # Blending factor
$\hat{{\bm{z}}}\leftarrow(1-\alpha){\bm{z}}_{c}+\alpha{\bm{z}}_{s}$;
  # Feature blending
$\hat{{\bm{i}}}_{\text{cs}}\leftarrow\ \mathcal{T}({\bm{i}}_{c},\hat{{\bm{z}}})$ $\beta\sim\mathcal{U}(\beta_{\text{min}},\beta_{\text{max}})$;
  # Interpolation factor
  # Stylized image
Algorithm 1 Style transfer augmentation block
@@
SASSL operates over minibatches, allowing efficient data preprocessing. Let ${\bm{I}}_{c}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ and ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{s}\times W_{s}}$ be content and style minibatches, respectively, comprised by $B$ images ${\bm{I}}_{c}^{(b)}$ and ${\bm{I}}_{s}^{(b)},b\in\{1,\dots,B\}$. Then, the stylized minibatch ${\bm{I}}_{\text{cs}}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying style transfer between a sample from the content batch and a sample from the style batch, given an arbitrary selection criterion. In what follows, we propose two alternatives for selecting style images to balance between augmentation diversity and efficiency.
@@
Diversifying style references. In contrast to traditional data augmentation, style transfer can leverage a second dataset to extract style references. This opens the possibility of selecting style images from different domains, diversifying the transformations applied to the pre-training dataset. SASSL relies on two approaches for sampling style references: external and in-batch stylization.
@@
External stylization consists on pre-computing representations of an arbitrary style dataset and sampling from them during pre-training. This allows controlling the styles to impose on the augmented views while reducing the computational overhead of Style Transfer. Under this configuration, the Style Transfer block receives a content minibatch along with a minibatch of pre-computed style representations extracted from an arbitrary style dataset, and generates a minibatch of stylized images.
@@
On the other hand, in-batch stylization uses the styles depicted in the content dataset itself by using other images of the content minibatch as style references. This is of particular interest for large-scale pre-training datasets covering multiple image categories and thus textures (e.g. ImageNet). So, enabling the use of a single dataset for both pre-training and stylization is a valid alternative.
@@
Let ${\bm{I}}_{c}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ be a pre-training minibatch comprised by $B$ images and ${\bm{I}}_{c}^{(b)}\in\mathbb{R}^{C\times H_{c}\times W_{c}},b\in\{1,\dots,B\}$ the image in the $b$-th index. Then, samples from the same minibatch can be used as style references by associating pairs of images in a circular fashion. More precisely, a style minibatch ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying a circular shift on the content minibatch indices
@@
where mod denotes the modulo operation and $b_{0}$ is an arbitrary offset. Figure 3 shows ImageNet samples stylized using pre-computed features from Painter by Numbers via external stylization, as well as using other ImageNet samples taken from the content minibatch via in-batch stylization.
Figure 2: Feature blending and image interpolation. A fine-grained control over the final stylization is obtained by introducing interpolation factors $\alpha$ and $\beta$ to operate in the feature space and pixel domain, respectively.
Content Image Style (In-batch) Style (External) Stylized (In-batch) Stylized (External) Refer to caption
Figure 2: Feature blending and image interpolation. A fine-grained control over the final stylization is obtained by introducing interpolation factors $\alpha$ and $\beta$ to operate in the feature space and pixel domain, respectively.
Figure 3: Style transfer examples. Stylized images generated using style references from the same domain (in-batch) as well as from other domain (external). Stylization obtained using a blending factor $\alpha=0.5$.
@@ 5.1 Downstream Task Performance
We compare the downstream ImageNet classification accuracy of SSL models learned using SASSL against a MoCo v2 (Chen et al., 2020b) baseline learned using the default data augmentation. Note that MoCo v2 and SimCLR are extremely similar (same loss, backbone architecture, and default augmentations), and differ mainly by MoCo’s use of a momentum encoder.
@@
Pre-training settings. Our pre-training setup is similar to the canonical SSL setup used to pre-train SimCLR and BYOL. We use the same loss, architecture, optimizer, and learning rate schedule as MoCo v2 for fair comparison. We pre-train a ResNet-50 encoder on ImageNet for $1,000$ epochs using our proposed method. Once the backbone is pre-trained, to measure downstream accuracy, we add a linear classification head on top of the backbone and train in a supervised fashion on ImageNet.
@@
SASSL applies Style Transfer only to the left view during pre-training, i.e., no changes in augmentation are applied to the right view. Style transfer is applied with a probability $p=0.8$ using blending and interpolation factors drawn from a uniform distribution $\alpha,\beta\sim\mathcal{U}(0.1,0.3)$. We found that this modest application of style transfer best complimented the existing augmentations, avoiding overly-strong augmentations that can reduce performance (Han et al., 2022; Chen et al., 2020a).
@@
Results. Table 1 compares the downstream classification accuracy obtained by our SASSL augmentation approach on MoCo v2 using external stylization from the Painter by Numbers dataset (Kan, 2016). Results indicate our proposed augmentation improves downstream task performance by $2.09\%$ top-1 accuracy. This highlights the value of Style Transfer augmentation in self-supervised training, where downstream task performance significantly boosts by incorporating transformations that decouple content and style. We also report results with in-batch stylization in Section 5.3.
Please refer to Section A.10 for additional ImageNet results showing SASSL’s performance on larger models (ResNet-50 x4 pre-trained via MoCo v2) and different architectures (ViT-B$/16$ pre-trained via MoCo v3). We experimentally show that SASSL also improves downstream task performance under these scenarios, highlighting its potential across different model complexities and architectures.
Table 1: Downstream classification performance on ImageNet. Classification accuracy of MoCo v2 pre-trained using our proposed style transfer augmentation approach (SASSL). Mean and standard deviation reported over five randomly initialized models.
@@
MoCo v2 (default) $72.55\pm 0.67$ $91.19\pm 0.34$
SASSL $+$ MoCo v2 (Ours) $\bf 74.64\pm 0.43$ $\bf 91.68\pm 0.36$
Table 2: Transfer learning performance. Downstream top-1 classification accuracy of MoCo v2 pre-trained on ImageNet using SASSL. SASSL generates specialized representations that improve transfer learning performance, both in linear probing and fine-tuning. Mean and standard deviation reported over five randomly initialized models.
@@
ImageNet ImageNet (1%) iNat21 Retinopathy DTD
Style Dataset Linear Probing
None (Baseline) $72.55\pm 0.67$ $53.23\pm 0.45$ $41.33\pm 0.2$ $\bf 75.88\pm 0.12$ $72.68\pm 0.7$
ImageNet (Ours) $74.07\pm 0.46$ $56.87\pm 0.43$ $45.01\pm 0.04$ $75.75\pm 0.1$ $73.69\pm 1.22$
iNat21 (Ours) $74.28\pm 0.38$ $56.76\pm 0.23$ $44.70\pm 0.37$ $75.75\pm 0.17$ $72.75\pm 1.01$
Retinopathy (Ours) $74.02\pm 0.61$ $\bf 56.99\pm 0.26$ $44.9\pm 0.16$ $75.78\pm 0.08$ $73.73\pm 0.57$
DTD (Ours) $74.32\pm 0.37$ $56.77\pm 0.36$ $\bf 45.08\pm 0.31$ $75.76\pm 0.11$ $\bf 74.41\pm 1.39$
PBN (Ours) $\bf 74.64\pm 0.43$ $56.9\pm 0.18$ $45.02\pm 0.14$ $75.79\pm 0.07$ $72.77\pm 0.77$
@@
None (Baseline) $74.89\pm 0.67$ $51.61\pm 0.13$ $77.92\pm 0.14$ $78.89\pm 0.2$ $71.54\pm 0.43$
ImageNet (Ours) $75.52\pm 0.23$ $51.74\pm 0.14$ $79.21\pm 0.07$ $79.64\pm 0.16$ $\bf 72.31\pm 1.85$
iNat21 (Ours) $\bf 75.58\pm 0.47$ $\bf 51.86\pm 0.3$ $79.19\pm 0.12$ $79.6\pm 0.23$ $71.35\pm 1.58$
Retinopathy (Ours) $75.52\pm 0.64$ $51.76\pm 0.26$ $79.23\pm 0.05$ $79.63\pm 0.13$ $72.07\pm 1.61$
DTD (Ours) $75.24\pm 0.65$ $51.73\pm 0.19$ $\bf 79.24\pm 0.08$ $\bf 79.7\pm 0.15$ $70.59\pm 1.42$
PBN (Ours) $75.05\pm 0.69$ $51.85\pm 0.16$ $79.2\pm 0.1$ $79.63\pm 0.13$ $71.35\pm 0.96$
5.2 Transfer Learning Performance
To understand robustness and generalization of representations learned using our approach, we study transfer learning performance across five tasks. By incorporating Style Transfer, we hypothesize that learned representations become invariant to changes in appearance (e.g. color and texture). This forces the feature extraction to rely exclusively on semantic attributes. As a result, the learned representations may become more robust to domain shifts, improving downstream task performance across datasets. We empirically show this by evaluating the transfer learning performance of representations trained using SASSL.
Downstream training settings. We compare the transfer learning performance of ResNet-50 pre-trained via MoCo v2 using SASSL against a vanilla MoCo v2 approach using default augmentations. Models are pre-trained on ImageNet and transferred to four target datasets: ImageNet 1% subset (Chen et al., 2020a), iNaturalist ‘21 (iNat21) (iNaturalist, 2021), Diabetic Retinopathy Detection (Retinopathy) (Kaggle & EyePacs, 2015), and Describable Textures Dataset (DTD) (Cimpoi et al., 2014). We also evaluate downstream performance on ImageNet, giving a total of five target datasets.
To have a clear idea of the effect of the style dataset in the SASSL pipeline, we pre-train five ResNet-50 backbones, each using a different style source. We cover the following style datasets: ImageNet, iNat21, Retinopathy, DTD, and Painter by Numbers (PBN). More precisely, we transfer five models, each pre-trained on a different style, to each of five target datasets. We also transfer a model pre-trained using the default augmentation pipeline. This leads to $30$ different transfer learning scenarios used to better understand the effect of various styles on different image domains.
Transfer learning performance is evaluated in terms of top-1 classification accuracy on linear probing and fine-tuning. All models were pre-trained as described in Section 5.1. We report mean and standard deviation across five model initializations. Please refer to Section A.4 for linear probing and fine-tuning training and testing settings.
Results. Table 2 shows the top-1 classification accuracy obtained via transfer learning. For linear probing, SASSL significantly improves the average performance on four out of five target datasets by up to $3.75\%$ top-1 classification accuracy. For Retinopathy, our augmentation method obtains on-par linear probing performance with respect to the default MoCo v2 model.
Similarly, for fine-tuning, all models trained via SASSL outperform the baseline. Results show the average top-1 classification accuracy improves by up to $1.32\%$. This suggests SASSL generalizes across datasets, spanning from texture (DTD) to the medical images (Retinopathy). Note that, for a fair comparison, we do not perform hyperparameter tuning on the default augmentations.
Table 3: Ablation Study. Linear probing accuracy for representations trained under different configurations. Mean and standard deviation reported over five randomly initialized models.
@@ 5.2 Transfer Learning Performance
MoCo v2 (default) -- -- 72.55±.67plus-or-minus72.55.6772.$5$67 $91.19\pm 0.34$
SASSL $+$ MoCo v2 (Ours) Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation $\beta=1$ ImageNet (in-batch) $70.87\pm 0.72$ $89.33\pm 0.34$
Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation: $\beta\in[0.1,0.3]$ ImageNet (in-batch) $74.07\pm 0.62$ $91.58\pm 0.18$
Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation: $\beta\in[0.1,0.3]$ PBN (external) $\bf 74.64\pm 0.43$ $\bf 91.68\pm 0.36$
5.3 Ablation Study
To better understand the impact of each proposed component, we conduct an ablation study over SASSL’s hyperparameters. Similar to previous experiments, we quantify our uncertainty by reporting the mean and variance obtained by five independent trials.

### v4

@@ 4 Proposed Method: SASSL
We provide a detailed description of our SSL augmentation technique. First, we break down SASSL’s key components and hyperparameters. Then, we tackle the problem of making the augmented images more diverse by utilizing style references from different domains in an efficient manner.
Output: ${\bm{i}}_{\text{cs}}$ ${\bm{z}}_{c}\leftarrow\mathcal{F}({\bm{i}}_{c})$ ;
Style representation of content image
${\bm{z}}_{s}\leftarrow\mathcal{F}({\bm{i}}_{s})$ ;
Style representation of style image
$\alpha\sim\mathcal{U}(\alpha_{\text{min}},\alpha_{\text{max}})$ ;
Blending factor
$\hat{{\bm{z}}}\leftarrow(1-\alpha){\bm{z}}_{c}+\alpha{\bm{z}}_{s}$ ;
Feature blending
$\hat{{\bm{i}}}_{\text{cs}}\leftarrow\ \mathcal{T}({\bm{i}}_{c},\hat{{\bm{z}}})$;
$\beta\sim\mathcal{U}(\beta_{\text{min}},\beta_{\text{max}})$ ;
Interpolation factor
Stylized image
Algorithm 1 Style transfer augmentation block
Content Image
Style (In-batch)
Style (External)
Stylized (In-batch)
Stylized (External)
[…]
figureStyle Transfer examples. Views generated using style references from the same domain (in-batch) as well as other domains (external). Stylization obtained using a blending factor $\alpha=0.5$.
@@
Style Transfer as data preprocessing. We incorporate Style Transfer to the default preprocessing pipeline of SSL methods. It is worth noting that SASSL is not specific to a particular SSL approach, and can be readily applied with different methods. Figure 1 shows an example of our augmentation pipeline, where Style Transfer is applied after random cropping. A raw input image ${\bm{i}}_{0}$ is cropped, producing a view that is taken as the content image ${\bm{i}}_{c}$. Given an arbitrary style image ${\bm{i}}_{s}$ (we discuss the choice of ${\bm{i}}_{s}$ below), the Style Transfer block generates a stylized image ${\bm{i}}_{\text{cs}}$ by imposing the texture of ${\bm{i}}_{s}$ over ${\bm{i}}_{c}$. Finally, the stylized image ${\bm{i}}_{\text{cs}}$ is passed to the remaining augmentation blocks to produce an augmented sample ${\bm{i}}_{\text{aug}}$.
@@
As discussed in recent work on SSL augmentation (Han et al., 2022; Chen et al., 2020a), adding a strong transformation to a self-supervised method tends to degrade performance. For this reason, it is crucial to control the amount of stylization imposed in the augmentation stage. We do so by introducing three hyperparameters: probability $p\in[0,1]$, which dictates whether an image is stylized or not, a blending factor $\alpha\in[0,1]$ to combine content ${\bm{z}}_{c}$ and style ${\bm{z}}_{s}$ representations, and an interpolation factor $\beta\in[0,1]$ to combine content ${\bm{i}}_{c}$ and stylized $\hat{{\bm{i}}}_{\text{cs}}$ images.
@@
Given style representations extracted from content and style images ${\bm{z}}_{c}=\mathcal{F}({\bm{i}}_{c})$ and ${\bm{z}}_{s}=\mathcal{F}({\bm{i}}_{s})$, respectively, we obtain an intermediate stylized image $\hat{{\bm{i}}}_{\text{cs}}$ by applying a convex combination based on blending factor $\alpha$.
$\hat{{\bm{i}}}_{\text{cs}}=$ $\ \mathcal{T}({\bm{i}}_{c},\hat{{\bm{z}}}),\quad\hat{{\bm{z}}}=\ (1-\alpha){\bm{z}}_{c}+\alpha{\bm{z}}_{s}$ (8)
@@
Then, the final stylization output is obtained via a convex combination between the intermediate stylized image $\hat{{\bm{i}}}_{cs}$ and the content image ${\bm{i}}_{c}$ based on interpolation factor $\beta$.
@@
Algorithm 1 describes our proposed Style Transfer data augmentation block. Figure 3.2 illustrates the effect of the feature blending and image interpolation operations, showcasing their importance to control the stylization effect without degrading the semantic attributes.
@@
SASSL operates over minibatches, allowing efficient data pre-processing. Let ${\bm{I}}_{c}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ and ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{s}\times W_{s}}$ be content and style minibatches, respectively, comprised by $B$ images ${\bm{I}}_{c}^{(b)}$ and ${\bm{I}}_{s}^{(b)},b\in\{1,\dots,B\}$. Then, the stylized minibatch ${\bm{I}}_{\text{cs}}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying Style Transfer between a sample from the content batch and a sample from the style batch, given an arbitrary selection criterion. We propose two alternatives for selecting style images to balance between augmentation diversity and efficiency.
@@
Diversifying style references. In contrast to traditional data augmentation, Style Transfer can leverage a second dataset to extract style references. This opens the possibility of selecting style images from different domains, diversifying the transformations applied to the pretraining dataset. SASSL relies on two approaches for sampling style references: external and in-batch stylization.
@@
External stylization consists on pre-computing representations of an arbitrary style dataset and sampling from them during pretraining. This allows controlling the styles to impose on the augmented views while reducing the computational overhead of Style Transfer. Under this configuration, the Style Transfer block receives a content minibatch along with a minibatch of pre-computed style representations extracted from an arbitrary style dataset, and generates a minibatch of stylized images.
@@
On the other hand, in-batch stylization uses the styles depicted in the content dataset itself by using other images of the content minibatch as style references. This is of particular interest for large-scale pretraining datasets covering multiple image categories and thus textures (e.g. ImageNet). So, enabling the use of a single dataset for both pretraining and stylization is a valid alternative.
@@
Following this, samples from the same minibatch can be used as style references by associating pairs of images in a circular fashion. More precisely, a style minibatch ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying a circular shift on the content minibatch indices
@@
where mod denotes the modulo operation and $b_{0}$ is an arbitrary offset. Figure 4 shows ImageNet samples stylized using pre-computed style representations from the Painter by Numbers dataset (Kan, 2016) via external stylization, as well as using other ImageNet samples taken from the content minibatch via in-batch stylization.
[…]
@@ 5.1 Downstream Task Performance
We evaluate the downstream ImageNet classification accuracy of SSL models pretrained via SASSL on the MoCo framework. We compare a MoCo v2 model pretrained with our data augmentation vs. a MoCo v2 baseline with default augmentation (Chen et al., 2020b). Note that MoCo v2 and SimCLR use the same loss, architecture, and augmentations (they differ by MoCo’s momentum encoding).
@@
Pretraining settings. Our pretraining setup is similar to the canonical SSL setup used to pretrain SimCLR and BYOL. We use the same loss, architecture, optimizer, and learning rate schedule as MoCo v2 for fair comparison. We pretrain a ResNet-50 encoder on ImageNet for $1,000$ epochs via SASSL. To measure downstream accuracy, we add a linear classification head on top of the pretrained backbone and train in a supervised fashion on ImageNet.
@@
SASSL pretraining applies Style Transfer only to the left view (no changes in augmentation are applied to the right view). It is applied with a probability $p=0.8$ using blending and interpolation factors drawn from a uniform distribution $\alpha,\beta\sim\mathcal{U}(0.1,0.3)$. We found that this modest stylization best complimented the existing augmentations, avoiding overly-strong transformations that can hinder performance (Han et al., 2022; Chen et al., 2020a).
@@
Results. Table 4 compares the downstream classification accuracy obtained by our SASSL augmentation approach on MoCo v2 using external stylization from the Painter by Numbers dataset. Results indicate our proposed augmentation improves downstream task performance by $2.09\%$ top-1 accuracy. This highlights the value of Style Transfer augmentation in self-supervised training, where downstream task performance significantly boosts by incorporating transformations that decouple content and style. We also report results with in-batch stylization in Section 5.5.
5.2 Transfer Learning Performance
To better understand the robustness and generalization of image representations learned using our proposed data augmentation approach, we evaluate its transfer learning performance across various tasks. By incorporating Style Transfer, we hypothesize that the learned representations become invariant to changes in appearance such as color and texture. This forces the feature extraction process to rely exclusively on semantic attributes. As a result, the learned representations may become more robust to domain shifts, improving downstream task performance across datasets. We empirically show this property by evaluating the transfer learning performance of image representations trained using SASSL under linear probing and fine-tuning scenarios.
Downstream settings. We compare the transfer learning accuracy of ResNet-50 pretrained via MoCo v2 using SASSL against a MoCo v2 baseline with default data augmentation. The evaluated models are pretrained on ImageNet and transferred to eleven target datasets: ImageNet-1% subset (Chen et al., 2020a), iNaturalist ‘21 (iNat21) (iNaturalist, 2021), Diabetic Retinopathy Detection (Retinopathy) (Kaggle & EyePacs, 2015), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Food101 (Bossard et al., 2014), CIFAR10/100 (Krizhevsky, 2009), SUN397 (Xiao et al., 2010), Cars (Krause et al., 2013), Caltech-101 (Fei-Fei et al., 2004), and Flowers (Nilsback & Zisserman, 2008).
To have a clear idea of the effect of the style dataset in SASSL’s pipeline, we pretrain five ResNet-50 backbones, each using a different style. We use ImageNet, iNat21, Retinopathy, DTD, and Painter by Numbers (PBN) as style datasets. More precisely, we transfer five models, each pretrained on a different style, to each of eleven target datasets. We also include ImageNet as target dataset to compare the effect of different styles on downstream task performance. This leads to $60$ transfer learning scenarios used to better understand the effect of various styles on different image domains.
Transfer learning is evaluated in terms of top-1 classification accuracy on linear probing and fine-tuning. All models were pretrained as described in Section 5.1. We report mean accuracy across five trials. Please refer to Section A.4 for full linear probing and fine-tuning training and testing settings.
@@ 5.2 Transfer Learning Performance
ImageNet ImageNet-1% iNat21 Retinopathy DTD Food101 CIFAR10 CIFAR100 SUN397 Cars Caltech-101 Flowers
Linear Probing
None (Default) $72.55$ $53.23$ $41.33$ $\bf 75.88$ $72.68$ $73.82$ $89.94$ $71.93$ $69.96$ $53.15$ $88.19$ $93.39$
ImageNet (Ours) $74.07$ $56.87$ $45.01$ $75.75$ $73.69$ $74.43$ $90.93$ $73.26$ $69.67$ $\bm{64.87}$ $89.3$ $95.27$
iNat21 (Ours) $74.28$ $56.76$ $44.70$ $75.75$ $72.75$ $74.3$ $\bm{91.04}$ $73.29$ $\bm{70.07}$ $63.96$ $\bm{89.89}$ $94.7$
Retinopathy (Ours) $74.02$ $\bf 56.99$ $44.9$ $75.78$ $73.73$ $74.53$ $90.8$ $73.3$ $69.63$ $64.06$ $89.17$ $94.94$
DTD (Ours) $74.32$ $56.77$ $\bf 45.08$ $75.76$ $\bf 74.41$ $\bm{74.88}$ $\bf{91.04}$ $\bm{73.41}$ $69.71$ $64.58$ $89.3$ $95.24$
PBN (Ours) $\bf 74.64$ $56.9$ $45.02$ $75.79$ $72.77$ $74.37$ $90.85$ $73.38$ $69.69$ $64.12$ $89.59$ $\bm{95.45}$
@@
None (Default) $74.89$ $51.61$ $77.92$ $78.89$ $71.54$ $87.25$ $96.91$ $83.4$ $74.25$ $83.63$ $89.27$ $95.75$
ImageNet (Ours) $75.52$ $51.74$ $79.21$ $79.64$ $\bf 72.31$ $87.48$ $97.0$ $83.21$ $73.89$ $\bm{90.33}$ $88.26$ $\bm{96.6}$
iNat21 (Ours) $\bf 75.58$ $\bf 51.86$ $79.19$ $79.6$ $71.35$ $87.4$ $\bm{97.05}$ $83.29$ $74.05$ $90.04$ $88.55$ $95.76$
Retinopathy (Ours) $75.52$ $51.76$ $79.23$ $79.63$ $72.07$ $87.39$ $96.97$ $\bm{83.68}$ $\bm{74.26}$ $89.96$ $88.44$ $96.34$
DTD (Ours) $75.24$ $51.73$ $\bf 79.24$ $\bf 79.7$ $70.59$ $\bm{87.66}$ $96.77$ $83.28$ $74.17$ $89.59$ $\bm{89.54}$ $95.59$
Style Dataset PBN (Ours) $75.05$ $51.85$ $79.2$ $79.63$ $71.35$ $87.56$ $96.97$ $83.36$ $74.18$ $89.75$ $88.97$ $95.77$
Results. Table 1 shows the top-1 classification accuracy obtained via transfer learning. For linear probing, SASSL significantly improves the average performance on eleven out of twelve target datasets by up to $10\%$ top-1 classification accuracy. For Retinopathy, SASSL obtains on-par linear probing accuracy to the default MoCo v2 model.
For fine-tuning, all models trained via SASSL outperform the baseline. Results show the average top-1 classification accuracy improves by up to $6\%$. This suggests SASSL generalizes across datasets, spanning from textures (DTD) to medical images (Retinopathy). Note that, for a fair comparison, we do not perform hyperparameter tuning. Interestingly, the relative performance obtained from different style datasets generally differs comparably to the measurement uncertainty, which is shown for these experiments in Section A.5 of the supplementary material. This suggests that the choice of style dataset is secondary in importance, while the main benefit comes from the use of SASSL itself.
@@
[…]
5.3 Few-shot Learning Performance
To further demonstrate the representation learning capabilities of the data augmentation imposed via SASSL, we conduct experiments on few-shot classification. We compare our ResNet-50 backbone pretrained via SASSL $+$ MoCo v2 against a MoCo v2 baseline in the context of one and ten-shot learning on ImageNet.
Table 5.2 shows the few-shot classification accuracy. Results reveal that SASSL boosts few-shot classification top-1 accuracy by over $1\%$ in both one and ten-shot learning. This aligns with our previous experiments, suggesting that SASSL promotes more general image representations.
5.4 Additional Downstream Performance Evaluation
Performance on other SSL methods. To assess SASSL’s broader impact, we evaluate its effectiveness on two other SSL methods, SimCLR and BYOL. We pretrain ResNet-50 backbones with each method, and then use linear probing on ImageNet to compare the quality of their learned representations. For each method, default pretraining and linear probing configurations are used. For SASSL, we employ its recommended hyperparameters ($\alpha,\beta\in[0.1,0.3]$, $p=0.8$) and PBN as style dataset.
Table 5.4 shows the accuracy attained by SimCLR and BYOL equipped with SASSL. Results show our proposed data augmentation technique boosts top-1 accuracy by approximately $1\%$ in both cases, highlighting its potential across multiple SSL techniques.
[…]
Performance on other representation models. We explore SASSL’s performance on models with varying complexity and architecture. For complexity, we employ ResNet-50 (x4), a scaled-up version of the previously evaluated ResNet-50 (from 24 to 375 million parameters). This allows us to probe how SASSL scales with increased model size. In terms of architecture, we employ ViT-B/16, a Transformer-based backbone with 86 million parameters and a distinct design compared to previous CNN models.
We pretrain and linearly probe a ResNet-50 (x4) representation model on ImageNet via MoCo v2. Pretraining and downstream settings follow our default configuration, as documented in the Sections A.3 and A.4. Similarly, we pretrain and linear probe a ViT-B/16 model on ImageNet via MoCo v3. In this case, SASSL employed a blending factor $\alpha$ uniformly sampled between $0.1$ and $0.5$.
Table 5.4 reports the downstream classification accuracy for ResNet-50 (x4) and ViT-B/16. ResNet-50 (x4) results show SASSL improves top-1 classification accuracy by $1.1\%$, mirroring its earlier improvement. Similarly, ViT-B/16 results show SASSL improves top-1 accuracy by $0.5\%$. These suggest that SASSL is not limited to CNN backbones, but can also be extended to ViTs. While this margin is currently smaller for ViTs, we emphasize that no hyperparameter tuning was employed in these experiments.
5.5 Ablation Studies
To shed light on how SASSL affects accuracy on ImageNet, we break down its components and assess individual contributions to downstream performance. We also study how aligning different layers in the stylization network $\mathcal{T}$ boosts accuracy. See Sections A.6 and A.7 for additional ablations and SASSL’s computational requirements.